What is the index of refraction of a refractive medium if the angle of incidence in air is 25 degrees and the angle of refraction is 15 degrees?

Highschool science 
Answer #1

(1) D

given,

the angle the incidence=60°

We know,

the angle of incidence= the angle of reflection

So, the angle of reflection= 60°

(2) Snell's low,

[N1Sine1 =N2Sine2]

N1=1.00

N2= 1.18

\Theta1=60°

\Theta2=?

so,

1.00×Sin60°= 1.18×Sin\Theta2

√3/2=1.18× Sin\Theta2

√3÷(2×1.18)= Sin\Theta2

1.732÷2.36= Sin\Theta2

0.733= Sin\Theta2

\Theta2=arcSin (0.733)

\Theta2 =47.13°

So, the angle of refraction=47.13°


answered by: ANURANJAN SARSAM